OpenStudy (2yoututu11):

The length of a rectangle is three times its width. If the perimeter is at most 112 centimeters, what is the greatest possible value for the width? Solve the problem. A. 42 cm B. 22.4 cm C. 14 cm D. 28 cm

OpenStudy (anonymous):

P=2(l+w) p=2(3w+w) p=8w 112=8w w=14...max. width would b 14

OpenStudy (campbell_st):

draw a picture and include the relevant information